OpenStudy (anonymous):

Suppose that y varies directly with x and inversely with z, and y = 18 when x = 15 and z = 5. Write the equation that models the relationship. Then find y when x = 21 and z = 7

OpenStudy (anonymous):

@Hero

hero (hero):

Hint: If y varies directly with x and inversely with z, then: yz = kx Input the given values to find k

OpenStudy (anonymous):

18(5) = k(15)?

hero (hero):

Yes, now solve for k: (18*5)/15 = k

OpenStudy (anonymous):

here are the multiple choice answers. y = 5z/x; 5/32 t = 5x/z;15 t = 6x/z;18 t = 6z/x;2

hero (hero):

Solving for k is only part of the problem. Find k, then get back to me.

OpenStudy (anonymous):

OK 18 * 5 IS 90/15 = k k = 6 @hero

hero (hero):

Okay, good. Now you have to use this formula again: yz = kx except this time use the fact that k = 6, x = 21, and z = 7 to find y

OpenStudy (anonymous):

y(7) = 21*6 7y = 126 y = 18

hero (hero):

GJ

OpenStudy (anonymous):

thanks so the answer t = 6x/z;18

OpenStudy (anonymous):

@hero

hero (hero):

I don't know where t comes from.

hero (hero):

If you have to choose one, then I suppose what you have chosen is correct.

OpenStudy (anonymous):

alright is there another way to solve this? that might give a different answer? @Hero

hero (hero):

There's only one way to solve it, and I have shown it to you. Your math course is being weird. Trust me. 99% of the world does it the way I've shown you (without the t)

hero (hero):

But don't worry. Your answer choice should be correct.

OpenStudy (anonymous):

alright thanks hun :P